# Efficient Dynamics Modeling in Interactive Environments with Koopman Theory

### v1

@@ 4 Experiments
While Koopman theory primarily focuses on state prediction, in reinforcement learning (RL) and planning tasks, the model needs to predict various other quantities as well, including rewards, value functions, and policies. These predictions are obtained by adding MLP heads on top of the representations. Hence, we give the dynamics model additional flexibility to learn useful latent representations. This is achieved by relaxing the consistency constraint of Eq. 11 using a loss coefficient that controls its contribution.
Section 4.1 presents our main empirical evaluation on long-range state and reward prediction, in interactive dynamics modeling tasks. Next, Section 4.2 presents the results of preliminary experiments on model-free RL and planning.
Table 1: Forward state and reward prediction error in Offline Reinforcement Learning environments. The reported number is the mean square error of the state and reward prediction over a horizon of 100 environment steps. The results are over 3 runs.
Offline Datasets State Prediction Error ($\downarrow$) Reward Prediction Error ($\downarrow$) MLP Transformer Koopman MLP Transformer Koopman hopper-expert-v2 0.500 $\pm$ 0.080 0.060 $\pm$ 0.007 0.007 $\pm$ 0.000 0.080 $\pm$ 0.005 0.030 $\pm$ 0.002 0.001 $\pm$ 0.000 hopper-medium-expert-v2 0.550 $\pm$ 0.120 0.063 $\pm$ 0.009 0.012 $\pm$ 0.004 0.260 $\pm$ 0.050 0.046 $\pm$ 0.008 0.002 $\pm$ 0.000 hopper-full-replay-v2 0.880 $\pm$ 0.150 0.240 $\pm$ 0.070 0.070 $\pm$ 0.004 0.400 $\pm$ 0.100 0.110 $\pm$ 0.040 0.012 $\pm$ 0.003 walker2d-expert-v2 0.300 $\pm$ 0.060 0.13 $\pm$ 0.060 0.090 $\pm$ 0.008 0.015 $\pm$ 0.003 0.004 $\pm$ 0.001 0.003 $\pm$ 0.000 walker2d-medium-expert-v2 1.320 $\pm$ 0.210 0.310 $\pm$ 0.070 0.220 $\pm$ 0.040 0.190 $\pm$ 0.030 0.032 $\pm$ 0.009 0.011 $\pm$ 0.002 walker2d-full-replay-v2 4.500 $\pm$ 2.100 3.300 $\pm$ 0.350 0.910 $\pm$ 0.050 0.520 $\pm$ 0.080 0.220 $\pm$ 0.020 0.090 $\pm$ 0.008 halfcheetah-expert-v2 2.860 $\pm$ 1.100 0.840 $\pm$ 0.160 0.300 $\pm$ 0.060 0.160 $\pm$0.050 0.070 $\pm$ 0.005 0.040 $\pm$ 0.006 halfcheetah-medium-expert-v2 0.590 $\pm$ 0.090 0.540 $\pm$ 0.071 0.568 $\pm$ 0.040 0.230 $\pm$ 0.050 0.120 $\pm$ 0.070 0.060 $\pm$ 0.005 halfcheetah-full-replay-v2 0.850 $\pm$ 0.200 0.930 $\pm$ 0.140 0.700 $\pm$ 0.040 0.170 $\pm$ 0.045 0.110 $\pm$ 0.052 0.091 $\pm$ 0.006
Dynamics Model Type Training sequence length 10 50 100 200 500 MLP 478.00 115.54 55.85 28.02 11.20 Transformer 418.00 120.20 73.34 34.48 14.92 Koopman 565.43 160.76 86.49 59.65 24.56
@@ 4.1 Long-Range Forward Dynamics Modeling with Control
We model the non-linear controlled dynamics of MuJoCo environments (mujoco, ) using our Koopman operator. We choose a standard MLP-based latent non-linear dynamics model for a baseline with two linear layers followed by a $\mathrm{ReLU}$ non-linearity. To make it easier to compare these two models, we use state embeddings of the same dimensionality using $g_{\theta}$, as described in Section 3.1. This embedding is then fed to the MLP-based latent dynamics model or the diagonal Koopman operator. This approach is widely used in RL for dynamics modeling pmlr-v162-hansen22a ; Silkchi2020LOOP ; schrittwieser2020mastering ; Schwarzer2020DataEfficientRL . To further compare our model with more expressive alternatives that can be trained in parallel, we design a causal Transformervaswani2017attention ; chen2021decision based dynamics model, which takes a masked sequence of state-actions and outputs representations that are used to predict next states and rewards. [^2] Typically all the states after the starting state are masked, but actions are fed to the model for all future time-steps. To ensure a fair comparison in terms of accuracy and runtime, we maintain an equal number of trainable parameters for all the aforementioned models. (see Appendix F for more details)
@@
For the forward dynamics modeling experiments, we use the D4RL fu2020d4rl dataset, which is a popular offline-RL environment. We select offline datasets of trajectories collected from three different popular Gym-MuJoCo control tasks: hopper, walker, and halfcheetah. These trajectories are obtained from three distinct quality levels of policies, offering a range of data representing various levels of expertise in the task: expert, medium-expert, and full replay. We divide the dataset of 1M samples into 80:20 splits for training and testing, respectively. To train the dynamics model, we randomly sample trajectories of length $\tau$ from the training data, where $\tau$ is the horizon specified during training. We test our learned dynamics model for a horizon length of 100 by randomly sampling 50,000 trajectories of length 100 from the test set. We use our learned dynamics model to predict the ground truth state sequence given the starting state and a sequence of actions.
4.1.1 Training Stability
@@
We can only train the MLP-based dynamics model for 10-time steps across all environments. Using longer sequences often results in exploding gradients and instability in training due to backpropagation through time sutskever2013training . The alternative of using ${\tanh}$ in MLP-based models can lead to vanishing gradients. In contrast, our diagonal Koopman operator handles long trajectories without encountering vanishing or exploding gradients, in agreement with our discussion in Section 3.2.1.
@@ 4.1.2 State and Reward Prediction
In Table 1, we evaluate our model’s accuracy in state and reward prediction over long horizons involving 100 environment steps. For this experiment, we add a reward predictor and jointly minimize the reward prediction loss and the state prediction loss. We set the weight of the consistency loss in the latent space to $0.001$. We see that for longer horizon prediction, our model is significantly better in predicting rewards and states accurately for hopper, walker, and halfcheetah. Furthermore, Table 2 empirically verifies that our proposed Koopman dynamics model is around twice as fast as an MLP or Transformer based dynamics model in learning from longer trajectories. The results also reveal the trend that our model’s relative speed-up over baselines grows with the length of the horizon.
Figure 2: Comparison of our Koopman-based dynamics model (with a horizon of 20) and an MLP-based dynamics model of vanilla TD-MPC pmlr-v162-hansen22a . The results are over 5 random seeds for each environment and are presented using the evaluation protocol from agarwal2021deep .
Figure 3: Comparison of vanilla SAC haarnoja2018soft and its integration with an MLP-based (SPR) and a Koopman-based dynamics model for incorporating self-predictive representations in the DeepMind Control Suite. The results are over 5 random seeds for each environment and are presented using the evaluation protocol from agarwal2021deep .
@@ 4.2 Koopman Dynamics Model for RL and Planning
We now present promising preliminary results in integrating the diagonal Koopman model into (I) model-based planning, and (II) model-free RL, where in the latter, the dynamics model is used for improved representation learning. While the same dynamics model can also be used for (III) model-based RL, we leave that direction for future work. Note that the distribution shift due to changing dynamics in this section hurts the evaluation of our diagonal Koopman model, and we hypothesize that the weaker results in this section are at least in part due to the non-IID setting.
@@
To assess the performance of our algorithm, we calculate the average episodic return at the conclusion of training and normalize it with the maximum score, i.e., 1000. However, due to significant variance over different runs, relying solely on mean values may not provide reliable metrics. To address this, agarwal2021deep suggests using bootstrapped confidence intervals (CI) with stratified sampling, which is particularly suitable for small sample sizes, such as the five runs per environment in our case. By utilizing bootstrapped CI, we obtain interval estimates indicating the range within which an algorithm’s aggregate performance is believed to fall and report the Interquartile Mean (IQM). The IQM represents the mean across the middle $50\%$ of the runs, providing a robust measure of central tendency. Furthermore, we calculate the Optimality Gap (OG), which quantifies the extent to which the algorithm falls short of achieving a normalized score of 1.0.  [^3] A smaller Optimality Gap indicates superior performance.
@@ 4.2.1 Model-Based Planning
We utilize TD-MPC pmlr-v162-hansen22a as the baseline for our model-based planning approach. TD-MPC combines the benefit of a model-free approach in long horizons by learning a value function while using an MLP-based “Task-Oriented Latent Dynamics” (TOLD) model for planning over shorter horizons. To assess the effectiveness of our proposed dynamics model, we replace TOLD with our diagonal Koopman operator. Subsequently, we trained the TD-MPC agent from scratch using this modified Koopman TD-MPC approach. See Section C.1 for details on this adaptation. To evaluate the performance of our variation, we conducted experiments in four distinct environments: Quadruped Run, Quadruped Walk, Cheetah Run, and Acrobot Swingup. We run experiments in the state space and compare them against vanilla TD-MPC.
Our Koopman dynamics model was trained with a horizon of 20-time steps. We observe that TD-MPC suffers from training instability, performance degradation and high variance when using 20-time steps. Consequently, we opted to use a horizon of 5-time steps for the vanilla approach. Fig. 2 suggests that the Koopman TD-MPC can be more efficient than the baseline. Since, due to high variance, this difference is not sufficiently pronounced, the reliability and stability of Koopman TD-MPC for a longer horizon is our main takeaway from this experiment.
@@ 4.2.2 Model-Free RL
Next, we apply our dynamics model in model-free RL, where we use an adaptation of Self-Predictive Representations (SPR; Schwarzer2020DataEfficientRL, ) as our baseline. This adaptation uses an MLP-based dynamics model rather than the original CNN used in SPR in order to eliminate the advantage of the original SPR for image inputs. SPR uses an auxiliary consistency loss similar to Eq. 11, as well as augmentations and other techniques from grill2020bootstrap to prevent representation collapse. It also improves the encoder’s representation quality by making it invariant to the transformations that are not relevant to the dynamics using augmentation; see also Srinivas2020CURLCU ; kostrikov2020image . To avoid overconstraining the latent space using the dynamics model, SPR uses projection heads. Similarly, we encode the Koopman observables using a projection layer above the representations used for Q-learning. We provide more details on the integration of the Koopman dynamics model into the SPR framework in Section C.2. The dynamics model is trained to predict representations up to $5$-time steps.
@@
We perform experiments on five different environments from the DeepMind Control (DMC) Suite tunyasuvunakool2020 ; tassa2018deepmind (Ball in Cup Catch, Cartpole Swingup, Cheetah Run, Finger Spin, and Walker Walk) in pixel space and compare it with an MLP-based dynamics model (SPR) and no dynamics model, i.e., SAC pmlr-v80-haarnoja18b as baselines. Figure 3 shows episode returns for these five environments. Koopman-based model outperforms the vanilla SAC and SAC plus MLP-based dynamics model.
Conclusion
@@
Dynamics modelling has a key role to play in sequential decision-making: at training time, having access to a model can lead to sample-efficient training, and at deployment, the model can be used for planning. Koopman theory is an attractive foundation for building such a dynamics model in Reinforcement Learning. The present article shows that models derived from this theoretical foundation can be made computationally efficient. We also demonstrate how this view gives insight into the stability of learning through gradient descent in these models. Empirically, we show the remarkable effectiveness of such an approach in long-range dynamics modelling and provide some preliminary yet promising results for model-based planning and model-free RL.
Limitations and Future Work
@@
Although our proposed Koopman-based dynamics model has produced promising results, our current treatment has certain limitations that motivate further investigation. Our current model is tailored for deterministic environments or tasks, overlooking stochastic dynamics modeling. We intend to expand our research in this direction, based on developments around stochastic Koopman theory mezic2005spectral , where Koopman observables become random variables, enabling uncertainty estimation in dynamics. Secondly, although our state prediction task yielded impressive results, we have identified areas for further growth, particularly in enhancing the applications of RL and planning. We hypothesize that the distribution shift during training and the objective change in training and evaluation hurt the performance of our Koopman dynamics model. We also plan to conduct a more comprehensive study involving a wider range of tasks and environments. Additionally, we aim to explore the compatibility of our model with different reinforcement learning algorithms to showcase its adaptability. Finally, we believe that the diagonal Koopman operator holds promise for model-based RL, where robust dynamics modeling over long horizons is precisely what is currently missing. We intend to investigate this direction, potentially unlocking new avenues for advancement.

### v4

@@ 4 Experiments
(a) hopper-v2
(b) walker2d-v2
(c) halfcheetah-v2
Figure 3: Forward state and reward prediction error in Offline Reinforcement Learning environments. We consider five dynamics modeling techniques and perform this prediction task over a horizon of 100 environment steps. The results are over 3 runs. Our Koopman-based method is competitive with the best performing GRU baseline while being 2×2$\times$2 × faster. See Appendix E for exact numerical values.
@@ 4.1 Long-Range Forward Dynamics Modeling with Control
Figure 4: Training speed in iterations/second ($\uparrow$) for the state prediction task using different dynamics model on halfcheetah-expert-v2. Each iteration consists of one gradient update of the entire model using a mini-batch of 256 in A100 GPU. See Table 3 for exact numerical values.
@@
We model the non-linear controlled dynamics of MuJoCo environments (Todorov et al., 2012) using our Koopman operator. We choose a standard MLP-based latent non-linear dynamics model with two linear layers followed by a $\mathrm{ReLU}$ non-linearity as one of our baselines. To make it easier to compare with Koopman dynamics model, we use state embeddings of the same dimensionality using $g_{\theta}$, which is an MLP, as described in Section 3.1. This embedding is then fed to the MLP-based latent dynamics model or the diagonal Koopman operator. This approach is widely used in RL for dynamics modeling (Sikchi et al., 2021, Hansen et al., 2022, Schrittwieser et al., 2020, Schwarzer et al., 2020). To further compare our model with more expressive alternatives that can be trained in parallel, we design a causal Transformer (Vaswani et al., 2017, Chen et al., 2021) based dynamics model, which takes a masked sequence of state-actions and outputs representations that are used to predict next states and rewards. [^3] All the states after the starting state are masked, but actions are fed to the model for all future time-steps. Following the success of Hafner et al. (2020), we also design a GRU-based dynamics model that also takes a masked sequence of state-actions to output representations for state and reward prediction. Finally, we use the same strategy to also design a DSSM-based Gu et al. (2022a) dynamics model. To ensure a fair comparison in terms of accuracy and runtime, we maintain an equal number of trainable parameters for all the aforementioned models. (see Appendix F for more details)
@@
For the forward dynamics modeling experiments, we use the D4RL Fu et al. (2020) dataset, which is a popular offline-RL environment. We select offline datasets of trajectories collected from three different popular Gym-MuJoCo control tasks: hopper, walker, and halfcheetah. These trajectories are obtained from three distinct quality levels of policies, offering a range of data representing various levels of expertise in the task: expert, medium-expert, and full replay. We divide the dataset of 1M samples into 80:20 splits for training and testing, respectively. To train the dynamics model, we randomly sample trajectories of length $\tau$ from the training data, where $\tau$ is the horizon specified during training. We test our learned dynamics model for a horizon length of 100 by randomly sampling 50,000 trajectories of length 100 from the test set. We use our learned dynamics model to predict the ground truth state sequence given the starting state and a sequence of actions.
Training stability
@@
We can only train the MLP-based dynamics model for 10-time steps across all environments. Using longer sequences often results in exploding gradients and instability in training due to backpropagation through time (Sutskever, 2013). The alternative of using ${\tanh}$ in MLP-based models can lead to vanishing gradients. In contrast, our diagonal Koopman operator handles long trajectories without encountering vanishing or exploding gradients, in agreement with our discussion in Section 3.2.1.
@@ 4.1.1 State and Reward Prediction
In Fig. 3, we evaluate our model’s accuracy in state and reward prediction over long horizons involving 100 environment steps. For this experiment, we add a reward predictor and jointly minimize the reward prediction loss and the state prediction loss. We set the weight of the consistency loss in the latent space to $0.001$. We see that for longer horizon prediction, our model is considerably better in predicting rewards and states accurately for hopper, walker, and halfcheetah in comparison to MLP, Transformer and DSSM based model while being competitive with GRU-based model. Furthermore, Fig. 4 empirically verifies that our proposed Koopman dynamics model is significantly faster than an MLP, GRU, DSSM and Transformer based dynamics model. The results also reveal the trend that our model’s relative speed-up over baselines grows with the length of the horizon.
Figure 5: Comparison of our Koopman-based dynamics model (with a horizon of 20) and an MLP-based dynamics model of vanilla TD-MPC (Hansen et al., 2022). The results are over 5 random seeds for each environment. Higher Mean &\&& IQM and lower Optimality Gap is better.
Figure 6: Comparison of vanilla SAC (Haarnoja et al., 2018) and its integration with an MLP-based (SPR) and a Koopman-based dynamics model for incorporating self-predictive representations in the DeepMind Control Suite. The results are over 5 random seeds for each environment. Higher Mean &\&& IQM and lower Optimality Gap is better.
@@ 4.2 Koopman Dynamics Model for RL and Planning
We now present promising results from integrating the diagonal Koopman model into two areas: (I) model-based planning, and (II) model-free reinforcement learning (RL). In the latter, the dynamics model enhances representation learning. Both approaches aim to solve continuous control tasks in a data-efficient manner. We provide a brief summary of these methods and their modifications with the Koopman dynamics model in Appendix C. While the same dynamics model can also be used for (III) model-based RL, we leave that direction for future work.
@@
To assess the performance of our algorithm, we calculate the average episodic return at the conclusion of training and normalize it with the maximum score, i.e., 1000. However, due to significant variance over different runs, relying solely on mean values may not provide reliable metrics. To address this, Agarwal et al. (2021) suggests using bootstrapped confidence intervals (CI) with stratified sampling, which is particularly suitable for small sample sizes, such as the five runs per environment in our case. By utilizing bootstrapped CI, we obtain interval estimates indicating the range within which an algorithm’s aggregate performance is believed to fall and report the Interquartile Mean (IQM). The IQM represents the mean across the middle $50\%$ of the runs, providing a robust measure of central tendency. Furthermore, we calculate the Optimality Gap (OG), which quantifies the extent to which the algorithm falls short of achieving a normalized score of 1.0.  [^4] A smaller Optimality Gap indicates superior performance.
@@ 4.2.1 Model-Based Planning
We utilize TD-MPC (Hansen et al., 2022) as the baseline for our model-based planning approach. TD-MPC combines the benefit of a model-free approach in long horizons by learning a value function while using an MLP-based “Task-Oriented Latent Dynamics” (TOLD) model for planning over shorter horizons. To assess the effectiveness of our proposed dynamics model, we replace TOLD with our diagonal Koopman operator. Subsequently, we trained the TD-MPC agent from scratch using this modified Koopman TD-MPC approach. See Section C.4 for details on this adaptation. To evaluate the performance of our variation, we conducted experiments in four distinct environments: Quadruped Run, Quadruped Walk, Cheetah Run, and Acrobot Swingup. We run experiments in the state space and compare them against vanilla TD-MPC.
Our Koopman dynamics model was trained with a horizon of 20-time steps. We observe that TD-MPC suffers from training instability, performance degradation and high variance when using 20-time steps. Consequently, we opted to use a horizon of 5-time steps for the vanilla approach. Fig. 5 suggests that the Koopman TD-MPC outperforms the MLP-based baseline while being more stable.
@@ 4.2.2 Model-Free RL
Current data-efficient model-free RL approaches like (SPR; Schwarzer et al., 2020) use dynamics modeling as an auxiliary task to improve representation learning. We apply our dynamics model to this model-free RL framework, where we use an adaptation of SPR as our baseline. This adaptation uses an MLP-based dynamics model rather than the original CNN used in SPR in order to eliminate the advantage of the original SPR for image inputs. SPR uses an auxiliary consistency loss similar to Eq. 6, as well as augmentations and other techniques from Grill et al. (2020) to prevent representation collapse. It also improves the encoder’s representation quality by making it invariant to the transformations that are not relevant to the dynamics using augmentation; see also Srinivas et al. (2020), Yarats et al. (2021). To avoid overconstraining the latent space using the dynamics model, SPR uses projection heads. Similarly, we encode the Koopman observables using a projection layer above the representations used for Q-learning. We provide more details on the integration of the Koopman dynamics model into the SPR framework in Section C.5.
@@
We perform experiments on five different environments from the DeepMind Control (DMC) Suite (Tunyasuvunakool et al., 2020, Tassa et al., 2018) (Ball in Cup Catch, Cartpole Swingup, Cheetah Run, Finger Spin, and Walker Walk) in pixel space, and compare it with an MLP-based dynamics model (SPR) and no dynamics model, i.e., SAC (Haarnoja et al., 2018) as baselines. Figure 6 demonstrates that for a horizon of 5 time steps Koopman-based model outperforms the SAC plus MLP-based dynamics model baseline while being more efficient.
5 Related Work
Diagonal state space models (DSSMs). DSSMs (Gu et al., 2022a, Gupta et al., 2022, Mehta et al., 2022) have been shown to be an efficient alternative to Transformers for long-range sequence modeling. DSSMs with certain initializations have been shown Gu et al. (2022a) to approximate convolution kernels derived for long-range memory using the HiPPO (Gu et al., 2020) theory. This explains why DSSMs can perform as well as structured state space models (S4) (Gu et al., 2022b), as pointed out by Gupta et al. (2022). Inspired by the success of gating in transformers (Hua et al., 2022), Mehta et al. (2022) introduced a gating mechanism in DSS to increase its efficiency further. However, a DSSM is a non-linear model used for sequence modeling and functions differently from our Koopman-based dynamics model, as explained in Appendix D.
Probabilistic latent spaces. Several early papers use approximate inference with linear and non-linear latent dynamics so as to provide probabilistic state representations (Haarnoja et al., 2016, Becker et al., 2019, Shaj et al., 2021, Hua et al., 2022, Fraccaro et al., 2017). A common theme in several of these articles is the use of Kalman filtering, using variational inference or MCMC, for estimating the posterior over the latents and mechanisms for disentangled latent representations. While these methods vary in their complexity, scalability, and representation power, similar to an LSTM or a GRU, they employ recurrence and, therefore, cannot be parallelized.
Koopman Theory for Control. The past work in this area has solely focused on learning the Koopman operator or discovering representations for Koopman spectrum for control (Shi & Meng, 2022, Watter et al., 2015). Korda & Mezić (2018) extends the Koopman operator to controlled dynamical systems by computing a finite-dimensional approximation of the operator. Furthermore, they demonstrate the use of these predictors for model predictive control (MPC). Han et al. (2020) tasks a data-driven approach to use neural networks to represent the Koopman operator in controlled dynamical system setting. Kaiser et al. (2021) introduces a method using Koopman eigenfunctions for effective linear reformulation and control of nonlinear dynamical systems. Song et al. (2021) proposes deep Koopman reinforcement learning that learns control tasks with a small amount of data. However, none of these works focus on the training speed and stability for predicting longer horizon.
6 Conclusion
@@
Dynamics modeling has a key role to play in sequential decision-making: at training time, having access to a model can lead to sample-efficient training, and at deployment, the model can be used for planning. Koopman theory is an attractive foundation for building such a dynamics model in Reinforcement Learning. The present article shows that models derived from this theoretical foundation can be made computationally efficient. We also demonstrate how this view gives insight into the stability of learning through gradient descent in these models. Empirically, we show the remarkable effectiveness of such an approach in long-range dynamics modeling and provide some preliminary yet promising results for model-based planning and model-free RL.
7 Limitations and Future Work
@@
Although our proposed Koopman-based dynamics model has produced promising results, our current treatment has certain limitations that motivate further investigation. Our current model is tailored for deterministic environments, overlooking stochastic dynamics modeling. We intend to expand our research in this direction, based on developments around stochastic Koopman theory (Mezić, 2005, Wanner & Mezic, 2022), where Koopman observables become random variables, enabling uncertainty estimation in dynamics. Secondly, although our state prediction task yielded impressive results, we have identified areas for further growth, particularly in its applications to RL and planning.
We hypothesize that the distribution shift during training and the objective change in training and evaluation hurt the performance of our Koopman dynamics model. We also plan to conduct a more comprehensive study involving a wider range of tasks and environments. Additionally, we aim to explore the compatibility of our model with different reinforcement learning algorithms to showcase its adaptability. Finally, we believe that the diagonal Koopman operator holds promise for model-based RL, where robust dynamics modeling over long horizons is precisely what is currently missing. We intend to investigate this direction, potentially unlocking new avenues for advancement.
Acknowledgements
This project was in part supported by CIFAR AI Chairs and NSERC Discovery grant. Computational resources are provided by Mila, ServiceNow Research and Digital Research Alliance (Compute Canada). The authors would like to thank Sebastien Paquet for his valuable feedback.